# Mobile Edge Intelligence for Large Language Models: A Contemporary Survey

### v1

@@ Abstract
On-device large language models (LLMs), referring to running LLMs on edge devices, have raised considerable interest owing to their superior privacy, reduced latency, and bandwidth saving. Nonetheless, the capabilities of on-device LLMs are intrinsically constrained by the limited capacity of edge devices compared to the much more powerful cloud centers. To bridge the gap between cloud-based and on-device AI, mobile edge intelligence (MEI) presents a viable solution to this problem by provisioning AI capabilities within the edge of mobile networks with improved privacy and latency relative to cloud computing. MEI sits between on-device AI and cloud-based AI, featuring wireless communications and more powerful computing resources than end devices. This article provides a contemporary survey on harnessing MEI for LLMs. We first cover the preliminaries of LLMs, starting with LLMs and MEI, followed by resource-efficient LLM techniques. We then illustrate several killer applications to demonstrate the need for deploying LLMs at the network edge and present an architectural overview of MEI for LLMs (MEI4LLM). Subsequently, we delve into various aspects of MEI4LLM, extensively covering edge LLM caching and delivery, edge LLM training, and edge LLM inference. Finally, we identify future research opportunities. We aim to inspire researchers in the field to leverage mobile edge computing to facilitate LLM deployment in close proximity to users, thereby unleashing the potential of LLMs across various privacy- and delay-sensitive applications.
@@ I-A Background
The recent advent of large language models (LLMs) has been a milestone in artificial intelligence (AI) technology to enable general-purpose intelligence. LLMs excel in various domains, i.e., not only in the task they are built for, i.e., generating text responses, but also in tasks such as multimodal content analysis, summarization, and generalization. For instance, the GPT-4 multimodal model accepts image and text inputs, which produces text outputs exhibiting human-level performance on various professional and academic benchmarks. Apart from these general-purpose models, sometimes called foundation models, LLMs can be fine-tuned to downstream tasks, catering to specific industries and application scenarios. For example, medical LLM, such as Med-PaLM M [1], was designed by Google to provide high-quality answers based on rich data modalities spanning text, imaging, genomics, and more. Google DeepMind also developed Robotics Transformer 2 (RT-2) [2], a vision-language-action AI model for controlling robots. The broad spectrum of use cases highlights the profound impact of LLMs on everyday lives.
Due to the associated computing, storage, and memory costs, the existing LLMs are mostly confined to cloud data centers for service provisioning. Regrettably, cloud-based LLM provisioning brings inherent drawbacks, including data privacy breaches, high bandwidth costs, and long service latency. Users must upload their data to utilize resources from cloud centers to access LLM services, which often results in significant communication delays. Moreover, uploading private data poses a serious risk to user privacy, especially in privacy-sensitive applications like smart health. Given these concerns, there is growing interest in on-device LLM deployment, which sparks a competitive race among major industry players. For instance, Google has launched Gemini Nano on Pixel 8 Pro smartphones with 1.8-billion and 3.25-billion parameters, respectively [3]. Qualcomm plans to launch Llama 2 support on Snapdragon-powered flagship smartphones and personal computers [4]. On-device LLM deployment enables local processing of sensitive personal data, such as end-to-end (E2E) encrypted messages and health data. It also provides low response time to delay-sensitive applications such as robot planning and autonomous driving. These salient advantages drive the continuing move of LLMs from cloud centers to mobile devices.
@@ I-B Motivation: From Cloud LLMs to On-device LLMs to MEI LLMs
Although on-device LLM is becoming a fast-growing field, the widespread deployment of on-device LLMs faces severe limitations. Specifically, the scarcity of computing, memory, and storage resources on edge devices substantially limits the scale of on-device LLM. On the one hand, the existing industrial efforts focus on sub-10B (10 billion parameters) LLMs due to the extensive resource requirements for the on-device deployment. For instance, Google’s Gemini Nano, which relies on 4-bit models with 1.8B and 3.25B parameters, can only support relatively “basic” functionalities like summarizing text, suggesting smart replies in context, and checking grammar [3]. However, as the desired functionalities become more complex, deploying larger-scale LLMs on the device becomes necessary, which can significantly increase the overhead for LLM on-device inference. On the other hand, on-device fine-tuning paves the way for personalized and context-aware AI, serving as a fundamental block for superior AI performance. However, the existing on-device LLM products do not incorporate on-device training (fine-tuning) functionalities due to the extensive training cost, which is usually much more intensive than AI inference.
To address the aforementioned dilemma, mobile edge computing offers a promising solution. The 6G mobile network aims to deliver low-latency AI inference and training services for a wide range of mobile devices based on edge computing systems by leveraging the network-endowed computing capabilities, say, on base stations. This leads to a paradigm called “mobile edge intelligence (MEI)”. MEI sits between on-device AI and cloud-based AI, featuring wireless communications and a modest scale of computing resources. In other words, it is more potent than edge devices, yet not as powerful as cloud centers. Due to the short distance between edge devices and edge servers, large-scale LLMs can be supported with lower service latency. Meanwhile, the 6G edge can continuously fine-tune LLMs in ever-evolving environments by exploiting the more capable memory, energy, and computing power on edge servers. As such, the 6G mobile edge is expected to be essential in pushing LLMs to edge devices.
@@ I-C Comparisons with Prior Surveys and Our Contributions
The deployment of LLMs is much more resource-intensive than conventional deep neural networks (DNNs), such as convolutional neural networks (CNNs), which is the main hurdle in bringing LLMs to the network edge. This survey paper aims to provide a contemporary survey of this converging trend, i.e., MEI and LLMs, primarily from the perspective of resource-efficient deployment, including storage efficiency, computing efficiency, and communication efficiency at the network edge. This paper differs from the prior survey papers on efficient LLM training/fine-tuning and inference, such as [14, 7, 9, 12, 13, 5, 15, 11]. These papers mostly focus on improving computing efficiency while overlooking the impact of communications on LLM training, inference, and caching and delivery, which is a significant bottleneck in mobile edge networks. This paper also differs from existing surveys/articles on LLM edge deployment, such as [16, 8, 10, 6]. These papers explore LLM-empowered AI service provisioning with cloud-edge synergy while not discussing resource-efficient deployments, such as parameter-efficient fine-tuning, split inference/learning, and efficient LLM caching and delivery and their interplay with wireless edge networks. Lastly, it is noted that this survey paper is fundamentally different from these papers on “LLMs for networks [17, 18], where the design objective is to employ LLMs to optimize edge networks rather than leveraging edge computing to support LLMs. The comparisons with some relevant surveys/papers are provided in Table I. The major contributions of this paper are summarized as follows.
@@
  • We present the application scenarios that drive the deployment of LLMs at the network edge. Although the use cases of LLMs have been extensively discussed in other places, we will emphasize the necessity or benefits of provisioning these applications at the mobile edge based on their service requirements.
@@
  • We provide the first comprehensive survey on how 6G edge networks facilitate LLM caching and delivery, training, and inference, including edge LLM caching and delivery, edge LLM training, and edge LLM inference. We will particularly concentrate on the resource-efficient deployment of LLMs to improve the storage, communication, and computing efficiency of LLMs at the network edge.
@@
  • We identify several crucial research directions for the integration of LLMs and mobile edge intelligence, including green edge AI and secure edge AI for LLMs.
@@
As the outline illustrated in Fig. 1, the survey is organized as follows. Section II presents an overview of LLMs and MEI, and Section III introduces the latest resource-efficient LLM techniques. Section IV illustrates four key applications that demonstrate the necessity of deploying LLMs at the network edge. In Section V, we present the MEI for LLM (MEI4LLM) framework, which supports the deployment of LLMs at the network edge. This framework consists of AI-native architecture, parameter-sharing LLM caching and delivery, distributed LLM training/fine-tuning, and distributed LLM inference. Sections VI, VII, and VIII delve into efficient techniques for edge LLM caching and delivery, edge LLM training, and edge LLM inference, respectively, considering storage efficiency, computing efficiency, and communication efficiency. Finally, we outline the roadmap for future research opportunities in Section IX and present our conclusions in Section X.
@@
II Preliminaries I: An Overview of LLMs and MEI
II-A Large Language Models
II-A1 Transformers
@@
Transformers work as follows. Unlike the recursive connections for short-term context and sequential processing in Recurrent Neural Networks (RNNs), Transformers employ self-attention mechanisms to comprehensively capture intricate dependencies between sequence elements to learn long-range relationships. The core of the Transformer architecture design lies in the encoder-decoder architecture, consisting of stacked layers with multi-head self-attention mechanisms. These mechanisms prioritize processing different elements in the input sequence, enhancing the model’s ability to effectively generate output tokens. Furthermore, each layer incorporates feed-forward networks (FFNs) and layer normalization. The encoder converts the input sequence into context-rich representations, whereas the decoder employs these representations to generate the output sequence, considering the input and previously generated tokens.
@@ II-A2 LLMs
The scalability of Transformers has fuelled the rise of LLMs. A variety of LLMs have been built and evolved based on Transformer architecture. Presently, major players in the AI industry are dedicated to crafting their LLMs and applying them across various domains. For instance, OpenAI has developed the highly-regarded chat LLM, GPT-3 [34], demonstrating exceptional performance across various NLP tasks, such as text generation and machine translation. Google has introduced the medical LLM, Med-PaLM [35], capable of offering expert-level medical guidance and diagnoses. Facebook proposed an innovative image classification LLM, DEiT [36], which integrates self-supervised learning with the Transformer architecture to achieve race-level image classification performance with limited annotated data. These LLMs are trained on extensive and varied datasets available on the Internet [37].
@@
LLM architectures can be classified into three categories: encoder-only LLMs, encoder-decoder LLMs, and decoder-only LLMs. Encoder-only LLMs, e.g. ALBERT [38], exclusively consist of encoder components and are typically based on advanced architectures such as Transformer [19]. The encoder is responsible for processing input sequences to generate contextualized representations for each token. Despite lacking a decoder to produce output sequences, encoder-only LLMs still exhibit exceptional performance in various NLP tasks such as text classification, sentence similarity computation, and language understanding due to their efficient feature extraction capabilities and adaptable representations. Encoder-decoder LLMs, exemplified by models like T5 [39], represent a pivotal advancement in the NLP domain, integrating both encoder and decoder components within their architectures. The encoder processes input sequences to generate contextualized representations, while the decoder utilizes these representations to generate output sequences, typically in a sequence-to-sequence manner. Encoder-decoder LLMs find widespread application in tasks such as machine translation, text summarization, and question answering, owing to their ability to capture complex linguistic structures and contextual dependencies. Decoder-only LLMs, epitomized by the well-known GPT series [34, 40], constitute a significant branch of LLMs. Decoder-only LLMs adopt autoregressive decoding, which is widely used in both decoder-only and encoder-decoder LLMs, to generate output sequences based on previous tokens in the sequence. This architectural design makes them particularly well-suited for tasks where the model generates text sequentially, such as language generation, text completion, and dialogue response generation.

### v2

@@ Abstract
On-device large language models (LLMs), referring to running LLMs on edge devices, have raised considerable interest since they are more cost-effective, latency-efficient, and privacy-preserving compared with the cloud paradigm. Nonetheless, the performance of on-device LLMs is intrinsically constrained by resource limitations on edge devices. Sitting between cloud and on-device AI, mobile edge intelligence (MEI) presents a viable solution by provisioning AI capabilities at the edge of mobile networks. This article provides a contemporary survey on harnessing MEI for LLMs. We begin by illustrating several killer applications to demonstrate the urgent need for deploying LLMs at the network edge. Next, we present the preliminaries of LLMs and MEI, followed by resource-efficient LLM techniques. We then present an architectural overview of MEI for LLMs (MEI4LLM), outlining its core components and how it supports the deployment of LLMs. Subsequently, we delve into various aspects of MEI4LLM, extensively covering edge LLM caching and delivery, edge LLM training, and edge LLM inference. Finally, we identify future research opportunities. We hope this article inspires researchers in the field to leverage mobile edge computing to facilitate LLM deployment, thereby unleashing the potential of LLMs across various privacy- and delay-sensitive applications.
@@ I-A Background
The recent advent of large language models (LLMs) has been a milestone in artificial intelligence (AI) technology to enable general-purpose intelligence. LLMs excel in various domains, i.e., not only in the task they are built for, i.e., generating text responses, but also in tasks such as multimodal content analysis, summarization, and generalization. For instance, the GPT-4 multimodal model accepts image and text inputs, which produces text outputs exhibiting human-level performance on various professional and academic benchmarks. Apart from these general-purpose models, sometimes called foundation models, LLMs can be fine-tuned to downstream tasks, catering to specific industries and application scenarios. For example, medical LLM, such as Med-PaLM M [1], was designed by Google to provide high-quality answers based on rich data modalities spanning text, imaging, genomics, and more. Google DeepMind also developed Robotics Transformer 2 (RT-2) [2], a vision-language-action AI model for controlling robots. The broad spectrum of use cases highlights the profound impact of LLMs on our everyday lives.
Due to extensive resource demands, most commercial LLMs are confined to cloud data centers for service provisioning. Regrettably, cloud-based LLM provisioning brings inherent drawbacks, including data privacy breaches, high bandwidth costs, and long service latency. Specifically, users must upload their data to cloud centers to access LLM services, resulting in significant communication delays. Moreover, uploading private data poses a serious risk to user privacy, especially in privacy-sensitive applications like smart health. During the transmissions and storage of data on cloud servers, unauthorized data access can expose sensitive personal information, such as medical records and biometric data, to malicious actors [3, 4, 5], leading to the misuse of data for malicious purposes and severely compromising user confidentiality. Given these concerns, there is a continuing development trend in on-device LLM deployment, sparking a competitive race among major industry players. For instance, Google has launched Gemini Nano on Pixel 8 Pro smartphones with 1.8-billion and 3.25-billion parameters, respectively [6]. Qualcomm plans to launch Llama 2 support on Snapdragon-powered flagship smartphones and personal computers [7]. On-device LLM deployment enables local processing of sensitive personal data and provides low response time, which is crucial for delay-sensitive applications such as robot planning and autonomous driving.
@@ I-B Motivation: From Cloud LLMs to On-device LLMs to MEI LLMs
Although on-device LLM is becoming a fast-growing field, the widespread deployment of on-device LLMs faces severe limitations. In particular, the scarcity of computing, memory, and storage resources on edge devices substantially limits the scale of on-device LLM. On the one hand, the existing industrial efforts focus on sub-10B (10 billion parameters) LLMs due to the extensive resource requirements for on-device deployment [8, 9]. For instance, Google’s Gemini Nano, which relies on 4-bit models with 1.8B and 3.25B parameters, can only support relatively “basic” functionalities like summarizing text, suggesting smart replies in context, and checking grammar [6]. However, as the desired functionalities become more complex, deploying larger-scale LLMs on the device becomes necessary. On the other hand, while on-device fine-tuning paves the way for personalized and context-aware AI, serving as a fundamental block for superior AI performance, the existing on-device LLM products generally do not incorporate on-device training (fine-tuning) functionalities due to the extensive training cost, which is much more resource-intensive than supporting AI inference alone.
To address the aforementioned dilemma, mobile edge computing offers a promising solution. The 6G mobile network aims to deliver low-latency AI inference and training services for a wide range of mobile devices by leveraging the network-endowed computing capabilities, say, on base stations. This leads to a paradigm called “mobile edge intelligence (MEI)”. Specifically, MEI sits between on-device AI and cloud-based AI, featuring a modest scale of computing resources located close to users, which is more capable than edge devices yet less powerful than cloud centers. Benefiting from the short distance between edge devices and edge servers, large-scale LLMs can be supported with lower service latency and bandwidth costs. Meanwhile, the 6G edge can continuously fine-tune LLMs for adapting to ever-evolving environments by exploiting the memory, energy, and computing power on edge servers, which is hard to achieve via edge devices alone. As such, the 6G mobile edge is expected to be essential in democratizing LLMs to edge devices, and providing comprehensive discussions and review on this trend is the focus of our survey paper.
@@ I-C Comparisons with Prior Surveys and Our Contributions
While MEI provisions powerful resources, supporting LLM services is still a non-trivial research task. The deployment of LLMs is much more resource-intensive than conventional deep neural networks (DNNs), such as convolutional neural networks (CNNs), which is the main hurdle in bringing LLMs to the network edge. Apart from communication-computing resource requirements [19, 20, 21], improving the resource and energy efficiency of LLMs is also crucial for successfully deploying them at the network edge. This survey paper aims to provide a contemporary survey of this converging trend, i.e., MEI and LLMs, primarily from the perspective of resource-efficient deployment of LLMs with MEI, including storage efficiency, computing efficiency, and communication efficiency at the network edge. This paper differs from the prior survey papers on efficient LLM training/fine-tuning and inference, such as [22, 12, 14, 17, 18, 10, 23, 16]. While these papers focus on improving computing efficiency, including efficient LLM compression methods [12], hardware optimization [14], efficient fine-tuning [18], and fast decoding [16], for efficient LLM training and inference, they overlook the impact of communications on LLM training, inference, and caching and delivery, which is a significant bottleneck in mobile edge networks. This paper also differs from existing surveys/articles on LLM edge deployment, such as [24, 13, 15, 11], which explore LLM-empowered AI service provisioning with cloud-edge synergy, including generative AI services [13], remote LLM inference [24], and LLM-driven sensing, digital twins, and task-oriented communications [25] in cloud-edge networks. Unfortunately, they do not discuss resource-efficient deployments, such as parameter-efficient fine-tuning, split inference/learning, efficient LLM caching and delivery, and their interplay with wireless edge networks. Lastly, it is noted that this survey paper is fundamentally different from these papers on “LLMs for networks” [26, 27, 28], where the design objective is to employ LLMs to optimize edge networks [26, 27] or telecommunication networks [28] rather than leveraging edge computing to support LLMs. Our paper, instead, explores the coordination of communication-computing resources in mobile edge networks to support LLM application provisioning, including caching, delivery, training, and inference of LLMs services at the network edge. The comparisons with some relevant surveys/papers are provided in Table I. The major contributions of this paper are summarized as follows.
@@
  • We present the application scenarios that motivate the deployment of LLMs at the network edge and the supporting MEI framework for LLM service provisioning. Although the use cases of LLMs have been extensively discussed in other places, we will emphasize the necessity or benefits of provisioning these applications at the mobile edge based on their service requirements, including latency, bandwidth, and privacy needs.
@@
  • We provide the first comprehensive survey on edge LLM caching and delivery, edge LLM training, and edge LLM inference within 6G edge networks. We will particularly concentrate on the resource-efficient deployment of LLMs to improve the storage, communication, and computing efficiency of LLMs at the network edge.
@@
  • We identify several crucial future research directions for the integration of LLMs and MEI, including green edge AI, secure edge AI, and quality-aware edge training for LLMs.
@@
As outlined in Fig. 1, the survey is organized as follows. Section II illustrates four key applications that demonstrate the necessity of deploying LLMs at the network edge. Section III presents an overview of LLMs and MEI, and Section IV introduces the state-of-the-art resource-efficient LLM techniques. In Section V, we present the MEI for LLM (MEI4LLM) framework, which supports the deployment of LLMs at the network edge. This framework consists of AI-native architecture, parameter-sharing LLM caching and delivery, distributed LLM training/fine-tuning, and distributed LLM inference. Sections VI, VII, and VIII delve into efficient techniques for edge LLM caching and delivery, edge LLM training, and edge LLM inference, respectively, considering storage efficiency, computing efficiency, and communication efficiency. Finally, we outline the roadmap for future research opportunities in Section IX and present our conclusions in Section X.
@@
II Application Scenarios
Although LLMs can be applied to a broad range of tasks, we focus on the application scenarios that motivate the deployment of LLMs at the network edge. As illustrated in Fig. 2, we will present four killer LLM-powered applications: mobile health, humanoid robots, virtual assistants, and autonomous driving. We focus on their service requirements, including latency, bandwidth, and privacy requirements, to underscore the need for deploying the LLMs at the network edge.
Latency reqs. $\leq$400 ms [29] 10-100 ms [30] 200 ms [31] 10 ms [30]
Lessons Learned: The stringent QoS requirements of LLM-empowered applications highlight the necessity of deploying LLMs in edge networks. First, applications demanding ultra-low latency cannot be supported by cloud computing due to the significant transmission delays over mobile backhaul and backbone networks. Second, ultra-high bandwidth costs resulting from large volumes of multimodal data not only raise the costs for both users and application developers (which need to pay the Internet service providers) but also result in network congestion. Third, with the sensitivity of the user input data, raw data sometimes must remain locally to prevent privacy leakage. These factors underscore the importance of deploying LLMs at the edge to provide timely, efficient, and privacy-preserving AI services to edge devices.
III Preliminaries I: An Overview of LLMs and MEI
As discussed in Section II, the emergence of LLMs has unlocked new avenues for enhancing various edge applications. To fully leverage MEI to support LLMs within edge networks, it is essential to comprehend the key concepts behind both LLMs and MEI. Towards this direction, Section III provides preliminaries of LLMs and MEI. Section IV reviews resource-efficient LLM techniques, which are indispensable for the edge deployment of LLMs. By laying out the foundational knowledge, these two sections prepare for the subsequent discussions on the integration of LLMs and MEI.
III-A Large Language Models
III-A1 The Transformer architecture
@@
A representative Transformer architecture is illustrated in Fig. 3. Unlike the recursive connections for short-term context and sequential processing in Recurrent Neural Networks (RNNs), Transformers employ self-attention mechanisms to comprehensively capture intricate dependencies between sequence elements to learn long-range relationships. The core of the Transformer architecture design lies in the encoder-decoder architecture, consisting of stacked layers with multi-head self-attention mechanisms. These mechanisms prioritize processing different elements in the input sequence, enhancing the model’s ability to generate output tokens effectively.
Figure 3: The Transformer architecture, adapted from [47].
Transformers typically operate as follows [11, 10, 47]. First, input data, such as words or sentences, are segmented into token sequences through a process called tokenization, using tokenizers like WordPiece [56, 57] or Byte-Pair Encoding [58, 51]. These transformed tokens are then converted into vectors by an embedding layer, which also incorporates position encoding to capture the input’s sequential information. Transformers then leverage the self-attention mechanism to evaluate the relationships between words within the inputs using their intermediate representations, i.e. query Q, key K, and value V, obtained by multiplying the input vectors with corresponding weight matrices. Afterward, layer normalization and feed-forward networks (FFNs) are applied. By using the self-attention mechanism and stacking multiple layers that integrate self-attention, layer normalization, and FFNs, the encoder converts the input sequence into context-rich representations, while the decoder employs these representations to generate the output sequence, considering the input and previously generated tokens. Finally, the output of the last Transformer layer is passed through a linear layer to generate the final result. Moreover, Transformers typically utilize autoregressive decoding to produce results. Each new token is predicted based on the input token sequence and then appended to the input sequence, forming a longer sequence for subsequent inference steps. To avoid redundant computation of previously processed tokens in earlier steps, a key-value (KV) cache can be employed, where the intermediate states, key, and value are stored after each inference step, allowing for more efficient processing.
@@ III-A2 Unimodal LLMs
LLMs mainly refer to advanced transformer-based language models consisting of billions of parameters. These models are extensively pre-trained on massive datasets using deep learning techniques [65, 10]. LLMs excel in tasks related to language comprehension, reasoning, and text generation, making them capable of solving complex tasks by producing coherent and contextually appropriate content. Notable examples include Meta’s LLaMA [7] and OpenAI’s GPT-4 [66]. Presently, major players in the AI industry are dedicated to crafting their LLMs and applying them across various domains. For instance, OpenAI has developed the highly-regarded chat LLM, GPT-3 [67], demonstrating exceptional performance across various NLP tasks, such as text generation and machine translation. Google has introduced the medical LLM, Med-PaLM [34], capable of offering expert-level medical guidance and diagnoses. Facebook proposed an innovative image classification LLM, DEiT [68], which integrates self-supervised learning with the Transformer architecture to achieve race-level image classification performance with limited annotated data. These LLMs are trained on extensive and varied datasets available on the Internet [69].
@@
A variety of LLMs have been built and evolved based on the Transformer architecture. LLM architectures can be classified into three categories: encoder-only LLMs, encoder-decoder LLMs, and decoder-only LLMs [28, 65, 11]. An overview of some popular LLMs with these three architectures is illustrated in Table II. Encoder-only LLMs, e.g. BERT [51] and ALBERT [70], exclusively consist of encoder components. The encoder is responsible for processing input sequences to generate contextualized representations for each token. Despite lacking a decoder to produce output sequences, encoder-only LLMs still exhibit exceptional performance in various NLP tasks such as text classification, sentence similarity computation, and language understanding due to their efficient feature extraction capabilities and adaptable representations. Encoder-decoder LLMs, exemplified by models like T5 [71], represent a pivotal advancement in the NLP domain, integrating both encoder and decoder components within their architectures. The encoder processes input sequences to generate contextualized representations, while the decoder utilizes these representations to generate output sequences, typically in a sequence-to-sequence manner. Encoder-decoder LLMs find widespread application in tasks such as machine translation, text summarization, and question answering, owing to their ability to capture complex linguistic structures and contextual dependencies. Decoder-only LLMs, epitomized by the well-known GPT series [67, 72], constitute a significant branch of LLMs, only consisting of a decoder network. This type of LLMs constitutes a significant branch of LLMs. An example of the decoder-only LLM architecture is illustrated in Fig. 4. Decoder-only LLMs adopt autoregressive decoding, which is widely used in both decoder-only and encoder-decoder LLMs, to generate output sequences based on previous tokens in the sequence. This architectural design makes them particularly well-suited for tasks where the model generates text sequentially, such as language generation, text completion, and dialogue response generation.
Figure 4: The decoder-only LLM architecture, which is adopted by GPT models [73].
TABLE II: An overview of some popular LLMs.
LLM arch. Model name Release Year Parameters
Encoder- only BERT [51] 2018 110-340M
ALBERT [70] 2019 12-235M
RoBERTa [74] 2019 355M
XLNet [75] 2019 110-340M
Decoder- only GPT-1 [73] 2018 117M
GPT-2 [76] 2019 1.5B
GPT-3 [77] 2020 175B
PaLM [78] 2022 8-540B
Gemini Nano-2 [79] 2023 3.25B
LLaMA [80] 2023 7-65B
GPT-4 [66] 2023 Close-sourced
Encoder- decoder T5 [71] 2019 60M-11B
BART [81] 2019 140M
mT5 [82] 2020 300M-13B
HuBERT [83] 2021 90M-1B
Claude 3 [84] 2024 Close-sourced